OpenStudy (anonymous):

Solve the following system of equations; 2x+y=3 Y=3-2x

OpenStudy (anonymous):

the second equation is the same as the first one... You can't solve two unknowns with one equation. You need another equation or there is no unique solution to the system

OpenStudy (texaschic101):

y = 3 - 2x -- (change to Ax + By = C) 2x + y = 3 same as other equation, which means the lines are coincident. They are the same. That means there is INFINITE SOLUTIONS

OpenStudy (anonymous):

Thank you